Carey Company had sales in 2016 of $1,768,500 on 65,500 units. Variable costs totaled $1,048,000, and fixed costs totaled $513,000. w raw material is available that will decrease the variable costs per unit by 20% (or S 3.20). However, to process the new raw material, fixed operating costs will increase by $91,000. Management feels that one-ha decline in the variable costs per unit should be passed on to customers in the form of a sales price reduction. The marketing department expects that this sales price reduction will result in a 5% increase in the nu of of units sold (a) Prepare a projected CVP income statement for 2017, assuming the changes have not been made. (Round per unit cost to 2 decimal places, e.g. 5.25 and all other answers to 0 decimal places, e.g. 1,225.) CAREY COMPANY CVP Income Statement Total Per Unit(b) Prepare a projected CVP income statement for 2017, assuming that changes are made as described. (Round per unit cost to 2 decimal places, e.g. 5.25 and all other answers to 0 decimal places, e.g. 1,225.) CAREY COMPANY CVP Income Statement 9 Total Per Unit
Answer #1

(a) Prepare a projected CVP income statement for 2017, assuming the changes have not been made,

Income Statement
Sales 1768500
Less : Variable Cost 1048000
Contribution Margin 720500
Less: Fixed Cost 513000
Net Operating Income 207500

(b) Assuming that changes are made as described.

Revised Variable cost per Unit = 1048000/65500 - 3.2 = 12.8

Revised Fixed Overhead =513000+91000=604000

Revised Sale Price per unit = 1768500/65500 -1/2*3.2 = 25.4

Revised Sale Quantity = 65500(1+5%) = 68775

Income statement
Sales ( 68775*25.40) 1746885
Less : Variable Cost ( 68775*12.8) 880320
Contribution Margin 866565
Less: Fixed Cost 604000
Net Operating Income 262565